OpenStudy (anonymous):

1/5x = 2x + 3 Please Help With Question.

OpenStudy (johnweldon1993):

Well you don't want 2 'x' numbers...so combine them... Subtract 2 from both sides of this equation. 1/5x - 2x = 3 1/5x - 2x = ?

OpenStudy (anonymous):

3?

OpenStudy (shamil98):

is this \[\frac{ 1 }{ 5 }x \] or \[\frac{ 1 }{ 5x }\]

OpenStudy (anonymous):